OpenStudy (jewlzme17):

Find the value x such that f(x)=0, if f(x)=(3x-4)/(5)

random231 (random231):

just solve (3x-4)/5=0

OpenStudy (jdoe0001):

\(\bf f(x)={\color{blue}{ y}}=0\qquad {\color{blue}{ f(x)}}=\cfrac{3x-4}{5}\qquad thus \\ \quad \\ {\color{blue}{ 0}}=\cfrac{3x-4}{5}\impliedby \textit{solve for "x"}\)

OpenStudy (jewlzme17):

I already know all of that I just don't remember how to solve (3x-4)/5

OpenStudy (freckles):

(3x-4)/5 there is nothing to solve here do you mean you don't remember how to solve (3x-4)/5=0?

OpenStudy (freckles):

the fraction is zero when the top is 0

OpenStudy (freckles):

that you only have to solve 3x-4=0

OpenStudy (jewlzme17):

so the zero cancels out the 5?

OpenStudy (dinamix):

(3x-4)=0 *5 3x=4 x=

OpenStudy (dinamix):

@jewlzme17 u should try little

OpenStudy (jewlzme17):

I know the answer it 4/3 its just I am having trouble understanding how. You guys are all saying different things and it's confusing. Like can someone just break it down barney style without being rude?

OpenStudy (freckles):

if 3x-4=0 then 3x has to be 4 another way to say this is 3x=4

OpenStudy (freckles):

can you solve for x now?

OpenStudy (freckles):

no one has said anything different

OpenStudy (freckles):

each of us has given you part of how to get there

OpenStudy (freckles):

@jdoe0001 showed you exactly what f(x)=0 meant @dinamix showed you how to get to 3x=4

OpenStudy (freckles):

oh and i forgot to add what I said @freckles said f(x)=0 when the numerator=0

OpenStudy (jewlzme17):

What I'm confused on is what you do with the denominator?

OpenStudy (freckles):

the bottom is always 5 it has no deciding factor when the fraction is 0

OpenStudy (freckles):

do you know that 0/anything is 0 (when anything isn't 0)

OpenStudy (freckles):

\[h(x)=\frac{f(x)}{g(x)}=0 \text{ when } f(x)=0 \text{ when } x \text{ is in the domain of } h \]

OpenStudy (freckles):

the only that gets to decide a fraction is 0 when the top is 0 on that function's domain of course

OpenStudy (jewlzme17):

OKay that makes so much more sense. Thank you

OpenStudy (freckles):

\[\frac{3x-4}{5}=0 \implies 3x-4=0\]

OpenStudy (freckles):

though you could have also decide to multiply both sides by 5 and that also works